Questionnare:

What are the measures of Angles a, b, and c? Show your work and explain your answers.

Questionnare:

1 attachment
Questionnare:

I remember doing these questions, but not with this many lines/angles...

Questionnare:

I see a right angle, but I don't see it associating with anything else.

carmelle:

Angle a is a vertical angle. So if the angle across from it is 30º, how many degrees is a? Once you know the value of a, you will use the trangle sum therom to find b. (Remember that there is a right angle) Triangle sum theorem: The sum of the measures of the interior angles of a triangle is 180° Angle c is a part of a straight angle, 180º, so subtract the value of the angle you already know to find the measrment of angle c.

carmelle:

measurement*

Questionnare:

I am not sure about Angle A.

Questionnare:

I know that Angle A is near the 30*

Questionnare:

Would it be 45*? Or perhaps 75*, similar to the other angle, C?

carmelle:

Vertical angles opposite each other. The measurement stays the same. So if the other angle is 30º, angle a is __º

Questionnare:

30*?

carmelle:

Correct

Questionnare:

Based on the Triangle Sum Theorem, I can, like you said, use angle A to find more info.

carmelle:

Yes.

Questionnare:

That means that I need to do: 30 + b + c = 180*?

carmelle:

You already know the measurment of two of your angles.

Questionnare:

30* + 75* + C* = 180*

Questionnare:

c* = 75?

carmelle:

no, look at where the arrow is pointing in the picture, that is a right angle. Right angles are how many degrees?

Questionnare:

90*

carmelle:

Yes, 90 + 30 + c = 180

Questionnare:

c = 60

carmelle:

correct, that means b = 60

Questionnare:

a = 30, b = 60, c = 60?

carmelle:

no, c is not 60. Remember what i said: "Angle c is a part of a straight angle, 180º, so subtract the value of the angle you already know to find the measrment of angle c."

Questionnare:

c = 90

carmelle:

How did you get that answer?

Questionnare:

I tried doing 30 + 60 + ? = 180, and got 90.

carmelle:

We are not using the triangle sum theorm for angle c. Subtract 180 from the angle given (its right next to angle c).

Questionnare:

180 - 75 = 105

carmelle:

Correct, c = 105

Questionnare:

oh k so...

Questionnare:

a = 30, b = 60, c = 105 no actually that doesn't add up to 180

Questionnare:

a = 15?

carmelle:

@questionnare wrote:
a = 30, b = 60, c = 105 no actually that doesn't add up to 180
there is no need to add to 180º anymore. (the triangle sum therom was only used to to find the measument of angle b)

Questionnare:

oh so a = 30 b = 60 c = 105

Questionnare:

okay i will type that down

carmelle:

@questionnare wrote:
oh so a = 30 b = 60 c = 105
Yes

Questionnare:

and then use the explanations

Questionnare:

thank you

carmelle:

You're welcome! :D